# Magnetic Structures Database from Symmetry-aided High-Throughput Calculations

### v3

@@ Abstract
Magnetic structures, which play a central role in determining their physical properties, are known for only very limited compounds. Traditional theoretical approaches to predicting magnetic structures predominantly rely on first-principles calculations. A key challenge of these methods is their requirement for initial magnetic configurations as inputs, which theoretically possess infinite possibilities. In this work, we introduce a strategy based on irreducible representation basis vectors that effectively narrows down the vast space of potential magnetic configurations to a finite set, typically comprising around 20 candidates per material. Despite this significant reduction, the compact input sets generated by our method already encompass the experimental magnetic structures for 253 out of 302 benchmark materials (83.8%) from the MAGNDATA database. These materials have propagation vectors $\mathbf{q}=0$ and unit cells containing up to 40 atoms, all within the Landau framework. Subsequent first-principles calculations correctly identify the magnetic structure in 198 of these cases. We further apply our highly efficient method to 7,520 stoichiometric transition-metal compounds with fewer than 20 atoms per unit cell in the Inorganic Crystal Structure Database, and establish a magnetic structure database containing 2,906 magnetic materials. To demonstrate its utility, we use this database for the systematic exploration of magnetic topological phases and altermagnets, identifying 1,070 and 392 candidate materials, respectively.
@@ I Introduction
In this work, we develop a magnetic-configuration strategy based on irreducible-representation (irrep) basis vectors. By construction, the resulting initial configuration space is extremely compact, typically comprising only about twenty candidates per material. Despite this drastic reduction, the generated candidate sets already contain the experimental magnetic structures for the majority of benchmark cases. Namely, for 253 out of 302 materials (83.8%) in the MAGNDATA database with propagation vector $\mathbf{q}=0$ and fewer than 40 atoms per unit cell, the experimental magnetic structures are explicitly included. Subsequent first-principles total-energy calculations then correctly identify the magnetic structure in 198 of these cases, corresponding to accuracy of 78.2%. Applying this approach to compounds with propagation vector $\mathbf{q}=0$ and fewer than 20 atoms per unit cell in ICSD Zagorac et al. (2019), we construct a database comprising over 2,900 magnetic materials. As the magnetic structure fundamentally governs all the physical properties of a material, our magnetic structure database provides a broader materials basis for systematic investigations of physical properties, such as spintronics, magnonics, magnetic topology, multiferroics, etc. Dagotto et al. (2001); Coey (2020); Bernevig et al. (2022); Eerenstein et al. (2006); Žutić et al. (2004); Han et al. (2020); Chumak et al. (2015); McClarty (2022); Broholm et al. (2020); Vojta (2018); Löhneysen et al. (2007); Šmejkal et al. (2022a, b); Ma et al. (2021). As an illustration of its capabilities, we present its application to the systematic identification of magnetic topological phases and altermagnets.
@@ II Method
Motivated by the excellent benchmark performance, we applied our framework to the Inorganic Crystal Structure Database (ICSD) Zagorac et al. (2019). Focusing on stoichiometric transition-metal compounds with fewer than 20 atoms per unit cell and excluding lanthanoids and actinoids, we restricted our analysis to magnetic structures with propagation vector $\mathbf{q}=0$. Within this well-defined scope, our irrep basis-vector strategy generates a compact set of trial magnetic configurations, typically <20<20 per material, enabling systematic high-throughput first-principles calculations. Out of 7,520 compounds examined, with a total of approximately 15,000 generated initial magnetic configurations, 4,614 materials are identified as non-magnetic. The remaining 2,906 magnetic materials, together with their magnetic structures, magCIF files, and associated metadata, have been compiled into a searchable database, available at https://magdatabase.nju.edu.cn.
@@ III.2 Magnetic topological semimetals
FM nodal-line semimetal Co3SnC: The Co3SnC (icsd_108129) adopts an anti-perovskite structure with the $Pm\bar{3}m$ (No. 221) space group Lin et al. (2014). Its magnetic ground state is a ferromagnet with magnetization oriented along the $[001]$ direction. The magnetic structure and the calculated electronic band structure of Co3SnC are shown in Fig. 3g and 3h. As shown in the Fig. 3i, two nodal rings appear on the high-symmetry planes $k_{z}=0$ and $k_{z}=\pi$. The two crossing bands carry +i+i and i-i eigenvalues of the mirror symmetry $m_{z}$, respectively, and are thus protected by $m_{z}$ symmetry. In addition, 16 symmetry-unprotected Weyl points are present in the Brillouin zone, with energies ranging from 0.1-0.1 to $0.15$ eV.
@@ V Outlook
The efficiency and scalability of our framework allow the construction of a comprehensive magnetic-structure database. In the present work, we focus on compounds with propagation vector $\mathbf{q}=0$ and fewer than 20 atoms per primitive cell as a starting point. Extensions to larger unit cells and to magnetic structures with $\mathbf{q}\neq 0$ are straightforward within our framework and are currently underway. Moreover, applying our strategy to two-dimensional systems is particularly timely and significant, given the rapidly growing interest in low-dimensional magnetism. Beyond the topological classification already established in this database, we are actively exploring additional magnetic-related physical properties, including multiferroic responses, anomalous Hall conductivity, transport coefficients, etc.

### v4

@@ Abstract
Magnetic structures, which play a central role in determining their physical properties, are known for only very limited compounds. Traditional theoretical approaches to predicting magnetic structures predominantly rely on first-principles calculations. A key challenge of these methods is their requirement for initial magnetic configurations as inputs, which theoretically possess infinite possibilities. In this work, we introduce a strategy based on irreducible representation basis vectors that effectively narrows down the vast space of potential magnetic configurations to a finite set, typically comprising around 20 candidates per material. Despite this significant reduction, the compact input sets generated by our method already encompass the experimental magnetic structures for 253 out of 302 benchmark materials (83.8%) from the MAGNDATA database. These materials have propagation vectors $\mathbf{q}=0$ and unit cells containing up to 40 atoms, all within the Landau framework. Subsequent first-principles calculations correctly identify the magnetic structure in 198 of these cases. We further apply our highly efficient method to 8,422 stoichiometric transition-metal compounds with fewer than 30 atoms per unit cell in the Inorganic Crystal Structure Database, and establish a magnetic structure database containing 2,906 magnetic materials. To demonstrate its utility, we use this database for the systematic exploration of magnetic topological phases and altermagnets, identifying 1,070 and 392 candidate materials, respectively.
@@ I Introduction
In this work, we develop a magnetic-configuration strategy based on irreducible-representation (irrep) basis vectors. By construction, the resulting initial configuration space is extremely compact, typically comprising only about twenty candidates per material. Despite this drastic reduction, the generated candidate sets already contain the experimental magnetic structures for the majority of benchmark cases. Namely, for 253 out of 302 materials (83.8%) in the MAGNDATA database with propagation vector $\mathbf{q}=0$ and fewer than 40 atoms per unit cell, the experimental magnetic structures are explicitly included. Subsequent first-principles total-energy calculations then correctly identify the magnetic structure in 198 of these cases, corresponding to accuracy of 78.2%. Applying this approach to compounds with propagation vector $\mathbf{q}=0$ and fewer than 30 atoms per unit cell in ICSD Zagorac et al. (2019), we construct a database comprising over 2,900 magnetic materials. As the magnetic structure fundamentally governs all the physical properties of a material, our magnetic structure database provides a broader materials basis for systematic investigations of physical properties, such as spintronics, magnonics, magnetic topology, multiferroics, etc. Dagotto et al. (2001); Coey (2020); Bernevig et al. (2022); Eerenstein et al. (2006); Žutić et al. (2004); Han et al. (2020); Chumak et al. (2015); McClarty (2022); Broholm et al. (2020); Vojta (2018); Löhneysen et al. (2007); Šmejkal et al. (2022a, b); Ma et al. (2021). As an illustration of its capabilities, we present its application to the systematic identification of magnetic topological phases and altermagnets.
@@ II Method
Motivated by the excellent benchmark performance, we applied our framework to the Inorganic Crystal Structure Database (ICSD) Zagorac et al. (2019). Focusing on stoichiometric transition-metal compounds with fewer than 30 atoms per unit cell and excluding lanthanoids and actinoids, we restricted our analysis to magnetic structures with propagation vector $\mathbf{q}=0$. Within this well-defined scope, our irrep basis-vector strategy generates a compact set of trial magnetic configurations, typically <20<20 per material, enabling systematic high-throughput first-principles calculations. Out of 8,422 compounds examined, with a total of approximately 170,000 generated initial magnetic configurations, 5,516 materials are identified as non-magnetic. The remaining 2,906 magnetic materials, together with their magnetic structures, magCIF files, and associated metadata, have been compiled into a searchable database, available at https://magdatabase.nju.edu.cn.
@@ III.2 Magnetic topological semimetals
FM nodal-line semimetal Co3SnC: The Co3SnC (icsd_108129) adopts an anti-perovskite structure with the $Pm\bar{3}m$ (No. 221) space group Lin et al. (2014). Its magnetic ground state is a ferromagnet with magnetization oriented along the $[001]$ direction. The magnetic structure and the calculated electronic band structure of Co3SnC are shown in Fig. 3g and 3h. As shown in the Fig. 3i, two nodal rings appear on the high-symmetry planes $k_{z}=0$ and $k_{z}=\pi$. The two crossing bands carry +i+i and i-i eigenvalues of the mirror symmetry $m_{z}$, respectively, and are thus protected by $m_{z}$ symmetry. In addition, 16 Weyl points are present in the Brillouin zone, with energies ranging from 0.1-0.1 to $0.15$ eV.
@@ V Outlook
The efficiency and scalability of our framework allow the construction of a comprehensive magnetic-structure database. In the present work, we focus on compounds with propagation vector $\mathbf{q}=0$ and fewer than 30 atoms per primitive cell as a starting point. Extensions to larger unit cells and to magnetic structures with $\mathbf{q}\neq 0$ are straightforward within our framework and are currently underway. Moreover, applying our strategy to two-dimensional systems is particularly timely and significant, given the rapidly growing interest in low-dimensional magnetism. Beyond the topological classification already established in this database, we are actively exploring additional magnetic-related physical properties, including multiferroic responses, anomalous Hall conductivity, transport coefficients, etc.